# BioNeMo Framework: A Modular, High-performance Library for AI Model Development in Drug Discovery

### v1

@@ 1 Introduction
The field of Artificial Intelligence (AI) has revolutionized computational biology and chemistry, culminating in the 2024 Nobel Prize in Chemistry for advancements in computational protein design and protein structure prediction. These breakthroughs have been driven by the rapid growth of high-throughput experimental biology [1], extensive sequence databases [2, 3], and significant progress in large-scale neural network architectures [4] leading to Large Language Models (LLMs). As such, protein Language Models (pLMs) [5, 6], trained on vast protein sequence databases [2], are now achieving unprecedented accuracy in tasks such as structure prediction [6], function prediction [7], and de novo protein design [8].
@@
With recent understanding of LLM scaling laws, pLMs have been growing in size to improve performance, with ESM3 featuring 98 billion parameters and requiring 1024 floating-point operations (FLOPS) to train [6, 9]. While improvements in training data quality may also help improve performance [10], we believe that frontier pLMs will continue to increase in size and training requirements. Beyond protein design, LLMs are also driving new techniques in chemical and material discovery [11, 12].
@@
AI-driven molecular research is driving a rapid increase in the need for specialized machine-learning frameworks tailored to drug discovery. While traditional tools like PyTorch[13], TensorFlow[14], and JAX[15] have been successful in advancing deep learning, these frameworks are not optimized for the complex biomolecular data and large-scale models required in drug discovery. Many of these frameworks are excellent for general-purpose deep learning but often require intricate adaptation for tasks such as protein structure prediction or chemical reaction modeling, resulting in trade-offs between flexibility and performance. Efficiently training very large models requires careful tuning of data and model-parallelism strategies and requires careful use of fused CUDA kernels to minimize memory usage and maximize training throughput [16][17]. High-performance data loaders are also a critical component of a pre-training pipeline, as ensuring that GPUs are fully utilized requires that pre-processing and memory access is done as efficiently as possible. Minimizing energy consumption and maximizing model performance are especially important as model training increases in scale. Further progress in these fields also requires that these techniques remain accessible to smaller research groups and domain scientists in chemical and biological research areas.
@@
In this paper, we present the BioNeMo Framework, a suite of open-source python packages designed to simplify the process of training high-performance AI models in the biomolecular and chemical space. These libraries build on top of NVIDIA NeMo[16] and Megatron-LM [17] to optimize training and inference throughput, and include high-performance data loaders for both protein sequence sampling and single-cell data. In this release, BioNeMo Framework includes the ESM-2 [6] and Geneformer [18] models, where we demonstrate an over 2x improvement in training throughput over PyTorch implementations and near-linear scaling in multi-device training up to 256 GPUs. We also discuss future directions for the BioNeMo Framework and how external contributors can extend the framework to accelerate additional families of models.
@@ 3.2 Benchmarks and Results
To assess the performance of the BioNeMo Framework, we compared training speeds for the ESM-2 model at various sizes and in different parallel configurations. First, we looked at single-device throughput for the 33-layer, 650M parameter variant. Benchmark speeds for the model as implemented in the Hugging Face Transformers library (4.44.0) were obtained using the Accelerate training library (0.33.0), using torch.compile’s default dynamo backend, running inside the NVIDIA Deep Graph Library (DGL) PyTorch container [^1] Access via nvcr.io/nvidia/pytorch:24.07-py3, PyTorch version 2.4.0). Training was performed on a single NVIDIA A100 GPU. Batches were increased until a CUDA out-of-memory (OOM) error occurred after a batch size of 16, with the resulting training throughput at each batch size shown in Figure 1. The BioNeMo Framework training throughputs were obtained on the same hardware with version 2.0 of the framework container. A maximum batch size of 46 was possible before an OOM error occurred, with an overall single-device training throughput 1.47x faster than that obtained with the Transformers library. BioNeMo Framework achieved a 59.2% model flops utilization (MFU) [19], while the model trained with the Accelerate library reached an MFU of 40.1%.
Next, we quantified how BioNeMo Framework performed in distributed training, at larger model sizes, and on H100 GPUs in addition to A100 GPUs. When training on 16 devices split across two compute nodes, the acceleration observed over the baseline training examples remained consistent. Training the 3B parameter variant of ESM-2 yields an MFU of 62.3% for BioNeMo Framework on A100 GPUs, which drops to 49.1% on the H100 system (a speedup of only 2.43x compared to the theoretical maximum of 3x). Hugging Face models were trained using DeepSpeed (0.15.1) with ZeRO level 1 using distributed data parallel [20].
@@ 3.5 Size-aware Batching
Many applications of neural network models require processing data of varying dimensions. Especially in the case of GNNs, the input data exhibit variations in their graph topologies, which are defined by the structure of their node sets and edge sets. In drug discovery and molecular modeling, graphs can represent the molecular bond structure of proteins or small molecules and the general spatial interactions among the atoms or higher-level chemical structures [21]. Instead of processing the data samples individually, processing them in batches is often more efficient for better hardware utilization via parallel data processing and for more stable gradient estimation during training, which can reduce training times. Perhaps the most obvious way of batching graphs is via padding, where a multi-dimensional array represents certain graph properties, with the leading dimension being the batch dimension, i.e., indexing the samples in the batch, while the remaining dimensions represent the inherent feature dimensions of the graph property in question. Padding is typically required in this scheme so that slicing along the batch dimension results in feature arrays of the same size. At the cost of extra memory space – and perhaps extra processing time due to the padding values – the padded array representation is convenient in indexing or slicing the samples, and it is usually machine learning framework friendly where a large number of arithmetic operations have been developed to work on dense arrays. A more memory-efficient approach to batching graph data leverages a graph adjacency matrix’s sparse representation to collate multiple graphs into a disjoint meta-graph, where sparse arrays represent the meta-graph itself. This approach typically concatenates samples’ array of a graph property, e.g., node features, into a single array and can discern the operations applied on different graphs in the batch via auxiliary data like an array that masks the individual graphs in the concatenation. No padding is required, and efficient algorithms have been developed to parallelize the sparse batch of samples subgraph sampling and message-passing techniques, among others. This approach is widely used in GNN frameworks such as PyG [22] and DGL [23]. While this approach is memory efficient and often saves unnecessary computation compared to the padding approach, it can become unintuitive as to what utility functions can be used for certain operations and what the time and space complexities are for those operations as a function of the input data size.
@@ 3.5.3 Benchmarks and results
Here we demonstrate the efficiency of size-aware batching in the GEOM-Drug data [24] by comparing three different data loading methods: the baseline data loading with static batch size, the MiDi data loader used by the Megalodon model, which implements an adaptive batch size [25], and our data loader using the size-aware batcher that takes the input samples from the buckets of the bucket batch sampler. We generate as many batches as 10 times the dataset size to simulate 10 training epochs using each method. We then examine these batches of data in terms of their data size or molecular size as well as the amount of padding involved. 4 shows the probability of samples of different sizes in the GEOM-Drug dataset. The MiDi and the baseline static size batching often undersample the moderate to large data for different reasons. The large samples are skipped in the baseline method due to OOM error. In the MiDi method, sampling is biased towards the most prominent sizes by the design of the method. The bucket size-aware batching gives a uniform distribution across a broad spectrum of data sizes, i.e., it does not introduce sampling bias.
@@ 4.1 Contributions from the Research Community
Flagship Pioneering is harnessing AI to drive greater innovation in the life sciences through their initiative Flagship Intelligence, which added protein inverse folding model RL-DIF[26] to the BioNeMo Framework. Unlike traditional protein folding models, which predict a structure from a given sequence, RL-DIF does the reverse: it generates diverse sequences that can fold into a specific target structure, allowing researchers to design new proteins that fit precise structural requirements and opening vast possibilities in drug discovery and biotechnology.

### v3

# references.bib

@@ 1 Introduction
The field of Artificial Intelligence (AI) has revolutionized computational biology and chemistry, culminating in the 2024 Nobel Prize in Chemistry for advancements in computational protein design and protein structure prediction. These breakthroughs have been driven by the rapid growth of high-throughput experimental biology [behjati2013next], extensive sequence databases [uniprot2023uniprot, lee2024petabase], and significant progress in large-scale neural network architectures [vaswani2017attention] leading to Large Language Models (LLMs). As such, protein Language Models (pLMs) [elnaggar2021prottrans, Lin2023-sb], trained on vast protein sequence databases [uniprot2023uniprot], are now achieving unprecedented accuracy in tasks such as structure prediction [Lin2023-sb], function prediction [bernhofer2021predictprotein], and de novo protein design [ferruz2023sequence].
@@
With recent understanding of LLM scaling laws, pLMs have been growing in size to improve performance, with ESM3 featuring 98 billion parameters and requiring 1024 floating-point operations (FLOPS) to train [Lin2023-sb, Kaplan2020-en]. While improvements in training data quality may also help improve performance [Fournier2024-fp], we believe that frontier pLMs will continue to increase in size and training requirements. Beyond protein design, LLMs are also driving new techniques in chemical and material discovery [Boiko2023-mj, Kononova2021-cq].
@@
AI-driven molecular research is driving a rapid increase in the need for specialized machine-learning frameworks tailored to drug discovery. While traditional tools like PyTorch[Paszke2019-ch], TensorFlow[tensorflow2015-whitepaper], and JAX[jax2018github] have been successful in advancing deep learning, these frameworks are not optimized for the complex biomolecular data and large-scale models required in drug discovery. Many of these frameworks are excellent for general-purpose deep learning but often require intricate adaptation for tasks such as protein structure prediction or chemical reaction modeling, resulting in trade-offs between flexibility and performance. Efficiently training very large models requires careful tuning of data and model-parallelism strategies and requires careful use of fused CUDA kernels to minimize memory usage and maximize training throughput [Kuchaiev2019-am][Narayanan2021-qc]. High-performance data loaders are also a critical component of a pre-training pipeline, as ensuring that GPUs are fully utilized requires that pre-processing and memory access is done as efficiently as possible. Minimizing energy consumption and maximizing model performance are especially important as model training increases in scale. Further progress in these fields also requires that these techniques remain accessible to smaller research groups and domain scientists in chemical and biological research areas.
@@
In this paper, we present the BioNeMo Framework, a suite of open-source python packages designed to simplify the process of training high-performance AI models in the biomolecular and chemical space. These libraries build on top of NVIDIA NeMo[Kuchaiev2019-am] and Megatron-LM [Narayanan2021-qc] to optimize training and inference throughput, and include high-performance data loaders for both protein sequence sampling and single-cell data. In this release, BioNeMo Framework includes the ESM-2 [Lin2023-sb] and Geneformer [Theodoris2023-np] models, where we demonstrate an over 2x improvement in training throughput over PyTorch implementations and near-linear scaling in multi-device training up to 256 GPUs. We also discuss future directions for the BioNeMo Framework and how external contributors can extend the framework to accelerate additional families of models.
@@ 3.2 Benchmarks and Results
To assess the performance of the BioNeMo Framework, we compared training speeds for the ESM-2 model at various sizes and in different parallel configurations. First, we looked at single-device throughput for the 33-layer, 650M parameter variant. Benchmark speeds for the model as implemented in the Hugging Face Transformers library (4.44.0) were obtained using the Accelerate training library (0.33.0), using torch.compile’s default dynamo backend, running inside the NVIDIA Deep Graph Library (DGL) PyTorch container [^1] Access via nvcr.io/nvidia/pytorch:24.07-py3, PyTorch version 2.4.0). Training was performed on a single NVIDIA A100 GPU. Batches were increased until a CUDA out-of-memory (OOM) error occurred after a batch size of 16, with the resulting training throughput at each batch size shown in Figure 1. The BioNeMo Framework training throughputs were obtained on the same hardware with version 2.0 of the framework container. A maximum batch size of 46 was possible before an OOM error occurred, with an overall single-device training throughput 1.47x faster than that obtained with the Transformers library. BioNeMo Framework achieved a 59.2% model flops utilization (MFU) [Chowdhery2022-hg], while the model trained with the Accelerate library reached an MFU of 40.1%.
Next, we quantified how BioNeMo Framework performed in distributed training, at larger model sizes, and on H100 GPUs in addition to A100 GPUs. When training on 16 devices split across two compute nodes, the acceleration observed over the baseline training examples remained consistent. Training the 3B parameter variant of ESM-2 yields an MFU of 62.3% for BioNeMo Framework on A100 GPUs, which drops to 49.1% on the H100 system (a speedup of only 2.43x compared to the theoretical maximum of 3x). Hugging Face models were trained using DeepSpeed (0.15.1) with ZeRO level 1 using distributed data parallel [Rajbhandari2019-wp].
@@ 3.5 Size-aware Batching
Many applications of neural network models require processing data of varying dimensions. Especially in the case of GNNs, the input data exhibit variations in their graph topologies, which are defined by the structure of their node sets and edge sets. In drug discovery and molecular modeling, graphs can represent the molecular bond structure of proteins or small molecules and the general spatial interactions among the atoms or higher-level chemical structures [David2020-gz]. Instead of processing the data samples individually, processing them in batches is often more efficient for better hardware utilization via parallel data processing and for more stable gradient estimation during training, which can reduce training times. Perhaps the most obvious way of batching graphs is via padding, where a multi-dimensional array represents certain graph properties, with the leading dimension being the batch dimension, i.e., indexing the samples in the batch, while the remaining dimensions represent the inherent feature dimensions of the graph property in question. Padding is typically required in this scheme so that slicing along the batch dimension results in feature arrays of the same size. At the cost of extra memory space – and perhaps extra processing time due to the padding values – the padded array representation is convenient in indexing or slicing the samples, and it is usually machine learning framework friendly where a large number of arithmetic operations have been developed to work on dense arrays. A more memory-efficient approach to batching graph data leverages a graph adjacency matrix’s sparse representation to collate multiple graphs into a disjoint meta-graph, where sparse arrays represent the meta-graph itself. This approach typically concatenates samples’ array of a graph property, e.g., node features, into a single array and can discern the operations applied on different graphs in the batch via auxiliary data like an array that masks the individual graphs in the concatenation. No padding is required, and efficient algorithms have been developed to parallelize the sparse batch of samples subgraph sampling and message-passing techniques, among others. This approach is widely used in GNN frameworks such as PyG [Fey2019-hu] and DGL [Wang2019-ch]. While this approach is memory efficient and often saves unnecessary computation compared to the padding approach, it can become unintuitive as to what utility functions can be used for certain operations and what the time and space complexities are for those operations as a function of the input data size.
@@ 3.5.3 Benchmarks and results
Here we demonstrate the efficiency of size-aware batching in the GEOM-Drug data [Axelrod2022-lo] by comparing three different data loading methods: the baseline data loading with static batch size, the MiDi data loader used by the Megalodon model, which implements an adaptive batch size [megaloden2024], and our data loader using the size-aware batcher that takes the input samples from the buckets of the bucket batch sampler. We generate as many batches as 10 times the dataset size to simulate 10 training epochs using each method. We then examine these batches of data in terms of their data size or molecular size as well as the amount of padding involved. 4 shows the probability of samples of different sizes in the GEOM-Drug dataset. The MiDi and the baseline static size batching often undersample the moderate to large data for different reasons. The large samples are skipped in the baseline method due to OOM error. In the MiDi method, sampling is biased towards the most prominent sizes by the design of the method. The bucket size-aware batching gives a uniform distribution across a broad spectrum of data sizes, i.e., it does not introduce sampling bias.
@@ 4.1 Contributions from the Research Community
Flagship Pioneering is harnessing AI to drive greater innovation in the life sciences through their initiative Flagship Intelligence, which added protein inverse folding model RL-DIF[ektefaie2024reinforcementlearningstructureconditionedcategorical] to the BioNeMo Framework. Unlike traditional protein folding models, which predict a structure from a given sequence, RL-DIF does the reverse: it generates diverse sequences that can fold into a specific target structure, allowing researchers to design new proteins that fit precise structural requirements and opening vast possibilities in drug discovery and biotechnology.